OpenStudy (anonymous):

Wally knows that in order to add or subtract rational expressions, he has to find the least common denominator first. Unfortunately, he can not remember how to do that. Using complete sentences, explain to Wally how to find least common denominators. Make sure you clearly explain any important items to consider.

OpenStudy (solomonzelman):

I'll give you an example, and you do the sentences part. \[\frac{3 }{20}+\frac{4 }{15}\]\[\frac{3 }{4 \times 5}+\frac{4 }{3 \times 5}\]\[\frac{3 }{\color{red} { 4 } \times 5}+\frac{4 }{\color{red} { 3 } \times 5}\]\[\frac{3\color{goldenrod} { \times 3 }}{4 \times 5\color{goldenrod} { \times 3 }} +\frac{3\color{goldenrod} { \times 4 } }{3 \times 5\color{goldenrod} { \times 4 }} \]\[\frac{3 \times 3 }{4 \times 5 \times 3} +\frac{3 \times 4 }{3 \times 5 \times 4} \]\[\frac{9 }{60} +\frac{12 }{60} \]\[\frac{9+12 }{60} \]\[\frac{21 }{60} \]\[\frac{7 \times \color{blue} { 3 } }{2 \times 5 \times 2 \times \color{blue} { 3 }} \]\[\frac{ 7} {2 \times 5 \times 2 }\]\[\frac{7}{20}\]

OpenStudy (solomonzelman):

I think this should be understandable.

OpenStudy (anonymous):

factor it?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

\[\frac{a}{b}\pm \frac{c}{d}=\frac{a d\pm b c}{b d} \] Wally does not need to know anything about least common denominators.

OpenStudy (solomonzelman):

That's a very good way of putting it @robtobey !

OpenStudy (solomonzelman):

I would say though that technically rally does need to know about Least Common Denominators.

OpenStudy (anonymous):

By the way,\[\frac{3}{20}+\frac{4}{15}=\frac{5}{12} \]

OpenStudy (solomonzelman):

Wally knows that in order to add or subtract rational expressions, he has to find the least common denominator first. Unfortunately, he can not remember how to do that.\(\color{red}{ Using~~~~complete~~~~sentences,~~~~explain~~~~to~~~~Wally~~~~how~~~~to~~~~find}\) \(\color{red}{ least~~~~common~~~~~~denominators. }\) Make sure you clearly explain any important items to consider. YOUR JOKE IS VERY RELATED THOUGH< CAN I GIVE YOU ANOTHER MEDAL?